In an alloy of 150 pounds of zinc and copper there are 100 pounds of copper. How much copper must be added so that the alloy may be 10% zinc.
Answer by ankor@dixie-net.com(22740)

In an alloy of 150 pounds of zinc and copper there are 100 pounds of copper.
How much copper must be added so that the alloy may be 10% zinc.
:
If there's 100 lb of copper there must be 50 lb zinc, initially
For a mixture to be 10% zinc, it has to be 90% copper
:
Write a percent copper equation
100 + x = .90(150+x)
100 + x = 135 + .9x
x -.9x = 135 - 100
.1x = 35
x = 35/.1
x = 350 lb of copper required
:
Check this; 350 + 150 = 500 lb resulting mixture of copper & zinc
Then .10(500) = 50 of zinc, which is the amt of zinc we have.
And .90(500) = 450 lb of copper